OpenStudy (emsxx3_):

what is the relationship between sine and cosine of complimentary angles and how do we know its true

OpenStudy (emsxx3_):

@Jadeishere

OpenStudy (jadeishere):

Can I help you in a few minutes, I'm helping two other people right now, won't take long! :)

OpenStudy (emsxx3_):

okay thank you

OpenStudy (jadeishere):

Okay, let's do this :)

OpenStudy (jadeishere):

So, do you know the SOHCAHTOA rule?

OpenStudy (emsxx3_):

no

OpenStudy (emsxx3_):

@Jadeishere

OpenStudy (emsxx3_):

the sine, cosine, and tangent of an angle. SOH stands for Sine equals Opposite over Hypotenuse. CAH stands for Cosine equals Adjacent over Hypotenuse. TOA stands for Tangent equals Opposite over Adjacent.

OpenStudy (emsxx3_):

is it that?

OpenStudy (jadeishere):

Yep xD It is that In relation to complimentary angles, there is a Complimentary Angle Theorem. Would you be willing to watch a video about it?

OpenStudy (jadeishere):

it explains it very well

OpenStudy (emsxx3_):

im kinda in a rush thats the thing

OpenStudy (emsxx3_):

@Jadeishere

OpenStudy (jadeishere):

Oh, well, the sin of an angle will be the same as the cos of the angles complimentary angle... DO you understand that? Kind of?

OpenStudy (emsxx3_):

kind of

OpenStudy (jadeishere):

They're co-functions, which is how we know it works. \[\cos(0) = \sin(90^{o} - 0)\] \[\sin (0) = \cos(90^o - 0)\]

OpenStudy (emsxx3_):

two cosine and sine are co functions ?

OpenStudy (emsxx3_):

@Jadeishere

OpenStudy (emsxx3_):

is what u wrote the answer ? @Jadeishere

OpenStudy (jadeishere):

So sorry I had to leave for a quick second! Yes :)

OpenStudy (emsxx3_):

so i would write that as my answers i add nothing ?

OpenStudy (emsxx3_):

thank you so much @Jadeishere

OpenStudy (jadeishere):

You can reword it if you'd like :) I love to help, tag me anytime :)